### v1

# Probing non-Gaussian correlations through entanglement generation in a many-body quantum system

Abstract
In understanding strongly correlated quantum systems, quantifying the non-Gaussian nature of interparticle correlations is invaluable. We show that, for a uniform quantum gas, there exists a natural connection between non-Gaussian correlations and the generation of momentum-space entanglement. Furthermore, this entanglement can be directly measured in an experiment using time-of-flight techniques. To prototype our method, we numerically study entanglement generation in a degenerate Bose gas following a quench to the unitary regime, where Gaussian and non-Gaussian correlations are generated sequentially in time.
Introduction.— Entanglement has attracted both great interest and considerable controversy since the early days of quantum mechanics, most famously expressed by the “EPR paradox” which posits that quantum mechanics is inconsistent with local realism Einstein et al. (1935); Reid et al. (2009). Today however, entanglement is well established as an empirical fact, and is regarded as a necessary resource for achieving quantum advantage in next-generation quantum technologies Acín et al. (2018); Preskill (2018). In addition to to these direct applications, the associated development of quantum information science has also sparked interest in using entanglement as a tool for studying many-body systems Amico et al. (2008); Horodecki et al. (2009); Hauke et al. (2016); Kunkel et al. (2022). For example, entanglement entropy has found use in condensed matter physics as an effective probe for detecting topological order and quantum critical behavior Osterloh et al. (2002); Vidal et al. (2003); Kitaev and Preskill (2006); Li and Haldane (2008); Jiang et al. (2012), and understanding thermalization in out-of-equilibrium quantum systems Eisert et al. (2015); Abanin et al. (2019); Chang et al. (2019); Mueller et al. (2022).
It is well known that generation of entanglement is intimately connected to the presence of non-Gaussian processes Walschaers (2021). For example, entanglement distillation of Gaussian states is only possible with non-Gaussian operations Eisert et al. (2002). This fact naturally leads to the question of whether entanglement can be used as a probe for quantifying the presence of non-Gaussian correlations in many-body quantum systems Strobel et al. (2014); Tenart et al. (2021); Bureik et al. (2025); Froland et al. (2024). Given the significant recent research interest in non-Gaussian processes, for instance in the generation of long-wavelength few-body correlations beyond the quasiparticle picture, such tools will be exceedingly valuable Armijo et al. (2010); Fletcher et al. (2017); Klauss et al. (2017); Schweigler et al. (2017); DeSalvo et al. (2019); Mestrom et al. (2020); Musolino et al. (2022); Patel et al. (2023); Zhang et al. (2023); Bureik et al. (2025). Particularly demanding are continuous variable systems in three dimensions, where current widespread theoretical approaches based on matrix product states and tensor networks become too demanding computationally, especially for studying non-equilibrium dynamics Haegeman et al. (2016); Kloss et al. (2018); Dutta et al. (2022); Ganahl et al. (2023).
Figure 1: Schematic representation of the four fundamental steps that generate intershell entanglement in the unitary Bose gas. Each panel shows the two-dimensional plane $k_{z}=0$ in the atomic (red) or molecular (green) space of momentum modes. Dots in the center represent the condensates, and shells represent spherically symmetric excitations at nonzero momentum.
In this Letter we show that, for a uniform quantum gas, the entanglement entropy between thin shells in momentum space can act as a witness for non-Gaussian processes Flynn et al. (2023). Furthermore, this entanglement entropy can be directly calculated by measuring density-density correlation functions, which are accessible through commonplace time-of-flight techniques Altman et al. (2004); Greiner et al. (2005); Cayla et al. (2018); Tenart et al. (2021). We present our formalism using a paradigmatic example of a three-dimensional quantum system containing strong non-Gaussian correlations, namely an ideal degenerate Bose gas quenched to the strongly interacting regime. Since the pioneering experimental works of Refs. Makotyn et al. (2014); Klauss et al. (2017); Eigen et al. (2017, 2018), the correlation dynamics in this system has received considerable theoretical interest Kira (2015a, b); Van Regemortel et al. (2018); Colussi et al. (2018a, b); Muñoz de las Heras et al. (2019); Colussi et al. (2020); Musolino et al. (2022); Ahmed-Braun et al. (2022); van de Kraats et al. (2024a). Initially, the system is dominated by Gaussian correlations, which induce the formation of a universal prethermal state where macroscopic observables look quasi-equilibrated Van Regemortel et al. (2018); Colussi et al. (2020). At later times, higher order non-Gaussian correlations develop, predominantly due to three-body scattering processes, which drive the system out of the prethermal stage and towards true thermalization, and generate off-diagonal long-range order in three-body correlations Musolino et al. (2022); van de Kraats et al. (2024a). We will show that this temporal separation of Gaussian and non-Gaussian dynamics becomes particularly natural when formulated in terms of momentum space entanglement.
Model.— We consider an ideal homogeneous $N$-particle Bose-Einstein Condensate (BEC) which is instantaneously quenched to the unitary regime, where the two-body s-wave scattering length $a\rightarrow\infty$ and interactions are as strong as allowed by quantum mechanics. The quench is initiated using a magnetic Feshbach resonance, which couples the atoms to a weakly bound diatomic molecular state Kohler et al. (2006); Chin et al. (2010). The subsequent quantum depletion of the condensate generates correlations between the excited momentum modes |$\absolutevalue*{\mathbf{k}}>0$, described by annihilation operators $\hat{a}_{\mathbf{k}}$. As pointed out in Ref. van de Kraats et al. (2024a), the post-quench correlation growth dynamics can be captured by the following two-point and three-point connected correlations functions, or cumulants,
$n_{\mathbf{k}}=\expectationvalue*{\hat{a}_{\mathbf{k}}^{\dagger}\hat{a}_{\mathbf{k}}},\ \kappa_{\mathbf{k}}=\expectationvalue*{\hat{a}_{\mathbf{k}}\hat{a}_{-\mathbf{k}}},\ R_{\mathbf{k},\mathbf{q}}=\expectationvalue*{\hat{a}_{\mathbf{k}}\hat{a}_{\mathbf{q}}\hat{a}_{-\mathbf{k}-\mathbf{q}}}.$ (1)
Here $n_{\mathbf{k}}$ is the excitation density, $\kappa_{\mathbf{k}}$ is the anomalous pairing density, and $R_{\mathbf{k},\mathbf{q}}$ the anomalous tripling density Colussi et al. (2018b, 2020). Eventually, higher order correlations are sequentially generated, restricting the range of times addressable with this truncated set Kira (2015a); Colussi et al. (2020). In the widely used Hartree-Fock-Bogoliubov (HFB) approximation Blaizot and Ripka (1985); Ahmed-Braun et al. (2022), three-body processes in the system are neglected by setting $R_{\mathbf{k},\mathbf{q}}=0$, which results in a fully integrable and Gaussian model of the dynamics, appropriate for describing early times following the quench. Upon including the non-Gaussian cumulant $R_{\mathbf{k},\mathbf{q}}$, the additional three-body scattering processes break integrability Van Regemortel et al. (2018), and significantly improve the match between model and experiment once the gas departs from the universal prethermal state van de Kraats et al. (2024a). Additionally, it was recently shown that $R_{\mathbf{k},\mathbf{q}}$ defines an order parameter for a novel Bose-Einstein condensate of correlated triples, which is particularly sensitive to the Efimov effect characteristic for strongly interacting three-body systems Musolino et al. (2022).
Entanglement in thin shells.— We follow the standard formulation of bipartite entanglement by dividing the system into regions $\mathcal{A}$ and $\mathcal{B}$, and quantify the entanglement between the two spaces with the Rényi entropy functional Rényi (1961); Horodecki et al. (2009),
$S_{\gamma}(\hat{\rho})=\frac{1}{1-\gamma}\ln\mathrm{Tr}\left[\hat{\rho}^{\gamma}\right]\quad\forall\gamma\in(0,1)\cup(1,\infty),$ (2)
where $\hat{\rho}=\mathrm{Tr}_{\mathcal{B}}\left[\outerproduct{\Psi}{\Psi}\right]$ is the reduced density matrix of the total many-body quantum state |$\ket*{\Psi}$. The structure of the relevant correlations in Eq. (1) suggests that a useful bipartition for probing non-Gaussianity is obtained by defining space $\mathcal{A}$ as containing all atomic momentum modes in a thin spherical shell with variable radius $k_{s}$ and width $\delta k_{s}$, thus placing all other atomic modes and all molecular modes in space $\mathcal{B}$ Flynn et al. (2023). This partition naturally separates the Gaussian correlation $\kappa_{\mathbf{k}}$, which correlates particles within a given shell, and the non-Gaussian correlation $R_{\mathbf{k},\mathbf{q}}$, which may correlate particles between different shells.
In Fig. 1 panels I-IV, we illustrate the application of the thin-shell partition to the dynamics of the quenched unitary Bose gas. Initially, association of atoms inside the BEC leads to the formation of a molecular condensate (I) Romans et al. (2004); Radzihovsky et al. (2004, 2008); Zhang et al. (2021); Sup . These molecules may then dissociate into correlated pairs (II), which must end up within a shell that defines a specific thin-shell bipartition. In the Gaussian model, this two-step excitation process is the only pathway for quantum depletion, implying that $\hat{\rho}$ is always a pure state with vanishing Rényi entropy. Note here that we assume the $U(1)$ symmetry breaking picture of Bose-Einstein condensation where the atomic and molecular condensates are described by classical fields, and consequently do not entangle with the excited modes Pitaevskii and Stringari (2016); Frérot and Roscilde (2015). If however, we account for the three-body correlations represented by $R_{\mathbf{k},\mathbf{q}}$, atoms within a shell can interact with the condensate to form excited molecules (III) van de Kraats et al. (2024a), whose subsequent decay can generate intershell entanglement (IV). This highlights the ergodic nature of this beyond-Gaussian model, and, most importantly for the purposes of this work, indicates how the entanglement entropy in space $\mathcal{A}$ provides a direct measure for the strength of non-Gaussian three-body correlations.
To quantify the entanglement entropy, we note that, within the coherent-state representation, $\hat{\rho}$ can be related directly to the cumulants in Eq. (1) Blaizot and Ripka (1985); Frérot and Roscilde (2015); Kira and Koch (2008); Serafini (2017); Sup . In this approach, the shell width $\delta k_{s}$ acts as a perturbative parameter, as has also been pointed out for interacting Fermi gases in Ref. Flynn et al. (2023). Specifically, retaining only the lowest order contributions in $\delta k_{s}$ gives the Gaussian state $\hat{\rho}=\hat{\rho}_{G}$, where,
$\hat{\rho}_{G}=\prod_{\mathbf{k}\in\mathcal{A}}\frac{1}{1+\tilde{n}_{\mathbf{k}}}e^{-\lambda_{\mathbf{k}}\hat{b}_{\mathbf{k}}^{\dagger}\hat{b}_{\mathbf{k}}},$ (3)
[…]
$\tilde{n}_{\mathbf{k}}=\frac{1}{2}\left(-1+\sqrt{(1+2n_{\mathbf{k}})^{2}-4\absolutevalue{\kappa_{\mathbf{k}}}^{2}}\right).$ (4)
[…]
$S_{\gamma}(\hat{\rho}_{G})=\frac{1}{\gamma-1}\sum_{\mathbf{k}\in\mathcal{A}}\ln\left[(1+\tilde{n}_{\mathbf{k}})^{\gamma}-\tilde{n}_{\mathbf{k}}^{\gamma}\right],$ (5)
which becomes an intensive quantity independent of $\delta k_{s}$ upon normalizing to the number of momentum modes $M_{\mathcal{A}}$ inside the shell, i.e. $s_{\gamma}(\hat{\rho}_{G})\equiv S_{\gamma}(\hat{\rho}_{G})/M_{\mathcal{A}}$.
It is important to understand that, even if non-Gaussian processes exist in the system, the thin-shell state in Eq. (3) is always Gaussian in form. The presence of non-Gaussian processes does however affect the spectrum $\lambda_{\mathbf{k}}$. Equation (5) suggests that in the Gaussian model, where the entanglement entropy vanishes, it must always hold that $\tilde{n}_{\mathbf{k}}=0$, meaning that the entanglement spectrum has an infinite gap. This condition is indeed enforced by the many-body equations of motion Sup , but attains a particularly elegant form in a phase space picture of the quantum state Kira (2015a). Within one half of the momentum shell ($\mathbf{k}>0$), we introduce the bidirectional basis $\hat{d}_{\mathbf{k}}=e^{-i\theta_{\mathbf{k}}}(\hat{a}_{\mathbf{k}}+\hat{a}_{-\mathbf{k}})/\sqrt{2}$ and fix the phase $\theta_{\mathbf{k}}$ such that $\expectationvalue*{\hat{d}_{\mathbf{k}}^{\dagger}\hat{d}_{\mathbf{k}}}=n_{\mathbf{k}}$ and $\expectationvalue*{\hat{d}_{\mathbf{k}}\hat{d}_{\mathbf{k}}}=\absolutevalue*{\kappa_{\mathbf{k}}}$ |. These operators define Hermitian field quadratures, $\hat{X}_{\mathbf{k}}=(\hat{d}_{\mathbf{k}}+\hat{d}_{\mathbf{k}}^{\dagger})/2$, and $\hat{Y}_{\mathbf{k}}=(\hat{d}_{\mathbf{k}}-\hat{d}_{\mathbf{k}}^{\dagger})/2i$, which obey the canonical commutation relation $[\hat{X}_{\mathbf{k}},\hat{Y}_{\mathbf{q}}]=i\delta_{\mathbf{k},\mathbf{q}}/2$. From here we can directly compute the standard deviations $\Delta X_{\mathbf{k}}$ and $\Delta Y_{\mathbf{k}}$, with associated Robertson uncertainty relation Robertson (1929),
$4\Delta X_{\mathbf{k}}\Delta Y_{\mathbf{k}}=(1+2n_{\mathbf{k}})^{2}-4\absolutevalue{\kappa_{\mathbf{k}}}^{2}\geq 1$ (6)
This inequality can also be derived from the positivity condition for the relative number squeezing parameter for $(\mathbf{k},-\mathbf{k})$ modes, as introduced in Refs. Roberts et al. (2002); Tenart et al. (2021). It is crucial to note however, that relative number squeezing indicates intrashell entanglement, which may be generated by Gaussian processes. Our thin-shell framework neglects this entanglement by construction, but rather focuses on the intershell entanglement, which arises out of the non-Gaussianity of correlations. Comparing with Eq. (4) we recognize that a vanishing eigenmode occupation, $\tilde{n}_{\mathbf{k}}=0$, corresponds with saturation of the uncertainty relation. Hence, $n_{\mathbf{k}}$ and |$\absolutevalue*{\kappa_{\mathbf{k}}}$ | together classify a thin-shell state as being: (1) forbidden ($\tilde{n}_{\mathbf{k}}<0$), (2) unentangled ($\tilde{n}_{\mathbf{k}}=0$) or (3) entangled ($\tilde{n}_{\mathbf{k}}>0$).
Before moving on to results, we want to emphasize that the formalism we have developed above presents a clear opportunity for an experiment. By measuring $n_{\mathbf{k}}$ and |$\absolutevalue{\kappa_{\mathbf{k}}}$ |, both of which are accessible through time-of-flight techniques Altman et al. (2004); Greiner et al. (2005); Cayla et al. (2018); Tenart et al. (2021), the entanglement spectrum and all derived quantities such as the entanglement entropy follow directly. As we have already argued, and will show in more detail in the remainder of this Letter, the entanglement characterized via this method then provides a direct probe of non-Gaussian processes in the system.
Results.— Having established the formalism for entropy generation in the unitary Bose gas, we will now characterize this entropy using a numerical simulation of the system. We employ the model developed in Ref. van de Kraats et al. (2024a), which formulates the many-body dynamics in terms of coupled nonlinear differential equations for a set of relevant atomic and molecular cumulants, including those presented in Eq. (1) (see Supplemental Material Sup for more details). In this model the interaction is characterized by the inverse scattering length $1/a$, which we fix at zero, and the intrinsic length $R_{*}$ Petrov (2004), which determines the molecular lifetime and characterizes the Feshbach resonance as broad ($R_{*}k_{n}\ll 1$) or narrow ($R_{*}k_{n}\gg 1$) Ho et al. (2012). As we discuss in more detail later, the intrinsic length can also be used to control the validity of the thin-shell approximation. Going forward, we will express all quantities in units of the Fermi momentum $k_{n}=(6\pi^{2}n)^{1/3}$ and Fermi time $t_{n}=2m/k_{n}^{2}$, where $n$ is the gas density and $m$ the atomic mass.
Figure 2: Trajectory through $(n_{\mathbf{k}},\absolutevalue{\kappa_{\mathbf{k}}}^{2})$ phase space for the Gaussian and beyond-Gaussian models in blue and red respectively, for $R_{*}k_{n}=0.4$ and $0\leq t/t_{n}\leq 3$. The gray region is forbidden by Eq. (6) and the white region is associated with intershell entanglement. Increasingly opague arrows show the direction of time. For visibility we rescale the axes of the four panels (a)-(d) by magnification factors $r=(1,2,10,50)$.
Figure 3: Growth of entanglement entropy in the unitary Bose gas. In Fig. (a), we plot the normalized second Rényi entropy $s_{2}(\hat{\rho}_{G})$ as a function of time and momentum in the beyond-Gaussian model, for $R_{*}k_{n}=0.4$. Blue dashed-dotted line highlights the characteristic shell momentum $k_{s}\xi=1$. In Fig. (b), we highlight the initial power law growth of $s_{2}(\hat{\rho}_{G})$, now at a set of resonance widths ranging from broad to narrow. Black dotted lines show the power law scaling $s_{2}(\hat{\rho}_{G})\sim(t/t_{*})^{6}$ expected to hold at early times $t\ll t_{*}$. To obtain the dimensionless proportionality constant, we fit this power law to the narrowest resonance in the third panel, and then use the same proportionality constant in the first and second panel. In Fig. (c), we plot the characteristic shell width $\Delta k_{s}^{\mathrm{max}}$ that determines the regime of validity for the thin-shell approximation, obtained with the correlation functions computed from our model and the experimental parameters of Ref. Eigen et al. (2018). Black horizontal line shows the bin size $\Delta k_{s}^{\mathrm{exp}}$ reported in Ref. Eigen et al. (2018) for time-of-flight measurements of the momentum space density.
In Fig. 2, we examine the trajectory of the many-body quantum state through the $(n_{\mathbf{k}},\absolutevalue{\kappa_{\mathbf{k}}}^{2})$ phase space, for Gaussian and beyond-Gaussian calculations. Following the quench at $t=0$, both the excitation density $n_{\mathbf{k}}$ and the pairing amplitude |$\absolutevalue*{\kappa_{\mathbf{k}}}$ | start to increase due to quantum depletion from the condensate. In the Gaussian model, |$\absolutevalue*{\kappa_{\mathbf{k}}}$ | grows in such a way that the state stays exactly on the boundary of the allowed and forbidden regions, as defined by Eq. (6). Hence, as predicted, this model does not generate entanglement. In contrast, in the beyond-Gaussian model the formation of correlated triples through the processes in the third and fourth panels of Fig. 1 slow the growth of the pairing amplitude |$\absolutevalue*{\kappa_{\mathbf{k}}}$ |, such that the state departs from the uncertainty boundary and travels into the entangled region of phase space.
Generally, $n_{\mathbf{k}}$ will keep growing following the quench until the system enters a quasi-equilibrated state, known as the prethermal state Van Regemortel et al. (2018); Colussi et al. (2020). Here, in the Gaussian model, $n_{\mathbf{k}}$ will plateau and start to oscillate coherently with a universal Bogoliubov frequency $\omega_{\mathbf{k}}$, analyzed in detail in Ref. Colussi et al. (2020). It defines the associated characteristic healing length $\xi k_{n}\approx\sqrt{2}$ and prethermalization timescale $\tau_{\mathbf{k}}\sim 1/\omega_{\mathbf{k}}$. For $k_{s}\xi\lesssim 1$, the characteristic excitations behave as long-wavelength phonons ($\omega_{\mathbf{k}}\sim k$), and the population oscillates slowly with respect to $t_{n}$. In the beyond-Gaussian model, we find that for these low-momentum modes the pairing amplitude itself reaches a plateau value, after which it stays approximately constant, such that the state gradually travels further into the entangled region Lat . For larger shell momenta $k_{s}\xi\gtrsim 1$ we enter the free-particle regime ($\omega_{\mathbf{k}}\sim k^{2}$), where the Gaussian model displays fast oscillations along the uncertainty boundary, seen clearly in Fig. 2. Here, in the beyond-Gaussian model, ergodic three-body correlations allow the system to break free from the prethermal state Van Regemortel et al. (2018); van de Kraats et al. (2024a), such that $n_{\mathbf{k}}$ increases beyond the maximum value attained in the Gaussian approximation. In this intermediate-time regime, the trajectory of the state through phase space becomes increasingly complicated, as seen particularly in the second and third panels of Fig. 2, however we find generally that the departure from the prethermal state is associated with a rapid damping of |$\absolutevalue*{\kappa_{\mathbf{k}}}$ |.
We now focus on the beyond-Gaussian model, and use Eq. (5) to calculate the second Rényi entropy ($\gamma=2$) as a function of both shell momentum and time following the quench, plotted in Fig. 3(a). As we show in the Supplemental Material Sup , our many-body model predicts that at early times $t\ll t_{*}$, $\tilde{n}_{\mathbf{k}}$ is independent of momentum and grows according to the power law $\tilde{n}_{\mathbf{k}}\sim(t/t_{*})^{6}$. Here $t_{*}=\sqrt{t_{n}\tau}$ is the geometric mean of the Fermi time $t_{n}$ and the molecular lifetime $\tau=mR_{*}/k_{n}$ Ahmed-Braun et al. (2022), and can be interpreted as the mean transition time for the processes shown in Fig. 1. It follows that, for $\gamma\geq 1$, the Rènyi entropy obeys the same early time scaling, i.e. $s_{\gamma}(\hat{\rho})\sim(t/t_{*})^{6}$, and indeed this prediction is confirmed by our numerics as shown for $\gamma=2$ in Fig. 3(b). At intermediate times, $t\gtrsim t_{*}$, this growth is gradually slowed down, and we find that while the entropy for $k_{s}\xi\lesssim 1$ still varies only weakly with $k_{s}$, the entropy for $k_{s}\xi\gtrsim 1$ now displays a rapid decrease with the shell momentum, see Fig. 3(a). This may be attributed to the universal suppression of excitations at large momenta due to the kinetic energy, which decreases the absolute value of the two point correlation functions $n_{\mathbf{k}}$ and $\kappa_{\mathbf{k}}$ Tan (2008a, b, c); Werner and Castin (2012a, b).
Validity of thin-shell approximation— In all results above we have considered just the lowest order contribution to the entropy with respect to the shell width $\delta k_{s}$. To examine the validity of this approximation, especially in regards to current experiments, we again employ the coherent-state representation of $\hat{\rho}$ to calculate the lowest order correction to Eq. (3), after which we obtain $\hat{\rho}=\hat{\rho}_{G}+\hat{\rho}^{\prime}$, where $\hat{\rho}^{\prime}$ is a non-Gaussian operator with vanishing trace Sup . Then, we appreciate that the difference between $\hat{\rho}$ and $\hat{\rho}_{G}$ can be quantified by the non-Gaussianity of the state $\hat{\rho}$, for which we use the following measure Genoni et al. (2007, 2008); Baek and Nha (2018); Park et al. (2021),
$\mathcal{N}(\hat{\rho}||\hat{\rho}_{G})=\frac{1}{2}\frac{\mathrm{Tr}\left[\left(\hat{\rho}-\hat{\rho}_{G}\right)^{2}\right]}{\mathrm{Tr}\left[\hat{\rho}^{2}\right]},$ (7)
which is a strictly positive function that vanishes if and only if $\hat{\rho}$ is indistinguishable from $\hat{\rho}_{G}$, in the sense that the Hilbert-Schmidt distance between the two states is zero. As we show in the Supplemental Material Sup , $\mathrm{Tr}\left[\hat{\rho}_{G}\hat{\rho}^{\prime}\right]=0$, such that the non-Gaussianity of $\hat{\rho}$ scales quadratically with $\hat{\rho}^{\prime}$. Specifically, we find,
$\mathcal{N}(\hat{\rho}||\hat{\rho}_{G})=\frac{1}{6}\sum_{\mathbf{k}\in\mathcal{A},\mathbf{q}\in\mathcal{A}}$ (8)
×[|$\times\left[\absolutevalue{R_{\mathbf{k},\mathbf{q}}}^{2}\mathcal{U}_{\mathbf{k},\mathbf{q}}-\left(R_{\mathbf{k},\mathbf{q}}^{*}\right)^{2}\mathcal{V}_{\mathbf{k},\mathbf{q}}-\left(R_{\mathbf{k},\mathbf{q}}\right)^{2}\mathcal{V}_{\mathbf{k},\mathbf{q}}^{*}\right].$
Here we have defined,
$\mathcal{U}_{\mathbf{k},\mathbf{q}}=\prod_{\mathbf{l}}\frac{\absolutevalue{u_{\mathbf{l}}}^{2}+\absolutevalue{v_{\mathbf{l}}}^{2}}{1+2\tilde{n}_{\mathbf{l}}},\quad\mathcal{V}_{\mathbf{k},\mathbf{q}}=4\prod_{\mathbf{l}}\frac{u_{\mathbf{l}}v_{\mathbf{l}}^{*}}{1+2\tilde{n}_{\mathbf{l}}},$ (9)
where $\mathbf{l}\equiv\left\{\mathbf{k},\mathbf{q},-\mathbf{k}-\mathbf{q}\right\}$. Due to the double momentum sum in Eq. (Probing non-Gaussian correlations through entanglement generation in a many-body quantum system), one finds that in the continuum limit $\mathcal{N}(\hat{\rho}||\hat{\rho}_{G})=\bar{\mathcal{N}}(\hat{\rho}||\hat{\rho}_{G})\delta k_{s}^{2}$, where $\bar{\mathcal{N}}(\hat{\rho}||\hat{\rho}_{G})$ is independent of the shell width. The thin-shell condition $\mathcal{N}(\hat{\rho}||\hat{\rho}_{G})\ll 1$ can then be recast as $\delta k_{s}\ll\Delta k_{s}^{\mathrm{max}}$, where $\Delta k_{s}^{\mathrm{max}}=\sqrt{1/\bar{\mathcal{N}}(\hat{\rho}||\hat{\rho}_{G})}$. Assuming that correlations become independent of the system volume in the thermodynamic limit, we can calculate $\Delta k_{s}^{\mathrm{max}}$ for a given experiment, and compare with the bin size $\Delta k_{s}^{\mathrm{exp}}$ for time-of-flight measurements, which determines the effective experimental shell width Ketterle et al. (1999). As an example, we use the quench experiment of Ref. Eigen et al. (2018), which reports a bin size $\Delta k_{s}^{\mathrm{exp}}\approx 0.02\ \mathrm{\mu m}^{-1}$ with respect to a total system volume $V=3\cdot 10^{4}\ \mathrm{\mu m}^{3}$ and Fermi momentum $k_{n}=6.7\ \mathrm{\mu m}^{-1}$. In Fig. 3(c) we show the resulting value of $\Delta k_{s}^{\mathrm{max}}$. Comparing with the bin size, we observe that the thin-shell approximation for this experiment is valid for the majority of momenta, but an improvement in resolution is desirable near the Fermi momentum at later times, where $R_{\mathbf{k},\mathbf{q}}$ is largest. We also observe that for narrower Feshbach resonances, the slower generation of beyond-Gaussian correlations as predicted in Refs. Ahmed-Braun et al. (2022); van de Kraats et al. (2024a), extends the validity of the thin-shell approximation, such that the width of the Feshbach resonance can be used as an additional control on the method. Hence, measuring thin-shell entanglement clearly presents both an opportunity and challenge for future experiments.
Its important to note that we have assumed that the initial many-body state is pure, which is only an approximation for ultracold experiments. Although theoretically more involved, notions of bipartite entanglement also exist for mixed states Amico et al. (2008), motivating an extension of our method in this direction. Going forward, our framework can be applied to study other quantum gases where non-Gaussian correlations are expected to be significant, including superfluid helium Dmowski et al. (2017); Tenart et al. (2021); Bureik et al. (2025), multicomponent Fermi gases Navon et al. (2010); Mukherjee et al. (2019); Ji et al. (2022); Schumacher et al. (2023), and Bose-Fermi mixtures Patel et al. (2023); Shen et al. (2024); van de Kraats et al. (2024b); Chuang et al. (2024). Our work also creates opportunities for further exploration of quantum resources in continuous variable systems, such as the entanglement spectrum Li and Haldane (2008) and quantum Fisher information Strobel et al. (2014); Hauke et al. (2016); Benavides-Riveros et al. (2024).
Acknowledgements.
We thank M. Wouters, R. Santos, J. Postema and R. de Keijzer for fruitful discussions. J.v.d.K. and S.J.J.M.F.K. acknowledge financial support from the Netherlands Organisation for Scientific research (NWO) under Grant No. 680.92.18.05, and from the Dutch Ministry of Economic Affairs and Climate Policy (EZK), as part of the Quantum Delta NL program. D.J.M.A.B. acknowledges funding by the Research Foundation–Flanders via a postdoctoral fellowship (Grant No. 1222425N).